OpenStudy (anonymous):

Find the polynomial function with the following roots: –2 of multiplicity 2, and 5

OpenStudy (anonymous):

is it -2^2=4 and -2^5=32?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

haha didnt think so

OpenStudy (anonymous):

first of all it is a polynomial that will look like \(x^3+bx^2+cx+d\) if one of the roots is 5 then one of the factors is \(x-5\)

OpenStudy (anonymous):

if one of the roots is \(-2\) then a factor is \(x+2\)

OpenStudy (anonymous):

and since you are told that \(-2\) is a root with "multiplicy" 2 that means you have a factor of \((x+2)^2\)

OpenStudy (anonymous):

meaning the answer is \[p(x)=(x+2)^2(x-5)\] but your teacher probably wants you to multiply this out

OpenStudy (anonymous):

do i foil it then?

OpenStudy (anonymous):

lol yeah "foil" aka multiply out

OpenStudy (anonymous):

\[(x+2)(x+2)(x-5)=(x^2+4x+4)(x-5)\] is a start

OpenStudy (anonymous):

um lol (x^2+4x)(x-1)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

wait, is it x^3-x^2-16x-20?

OpenStudy (phi):

you can use the distributive property A ( x + y) = Ax + A y where A can be complicated \[ (x^2+4x+4)(x-5) = (x^2+4x+4)x + (x^2+4x+4)(-5) \] now distribute again, and combine "like terms" your answer looks good but I did not check it carefully

OpenStudy (anonymous):

okay, thank you!!